OpenStudy (anonymous):

solve the system by substitution 1.) 6x-8y=-16 x+y=-5 2.) y=7x+14 y= 7x+14 Solve the system by elimination 3.) x-3y=14 -7x+9y=-14 4.) -7x-7y=-7 8x+8y=8

OpenStudy (anonymous):

The first step in substitution is to isolate the variable. for example one pick one of the variable to isolate. x=y-5

OpenStudy (anonymous):

Then you plug in that expression into the equation and solve 6(y-5) + 8y =16

OpenStudy (anonymous):

14y-30=16

OpenStudy (anonymous):

14y=46

OpenStudy (anonymous):

y=7/23

OpenStudy (anonymous):

next you plug that value into the original equation and find the other variable

OpenStudy (anonymous):

By this you are substituting the values

OpenStudy (anonymous):

Didn't see the negative

OpenStudy (anonymous):

(-6,1) is your answer